OpenStudy (anonymous):

i give medal Determine which system below will produce infinitely many solutions. A. -6x + 3y = 18 4x - 3y = 6 B. 2x + 4y = 24 6x + 12y = 36 C. 3x - y = 14 -9x + 3y = -42 D. 5x + 2y = 13 -x + 4y = -6

OpenStudy (anonymous):

c

OpenStudy (lyrae):

You have to check if one of the systems have equal equations. You do this by trying to multiply one equation in each system by some constant to make it equal to the other equation in the same system. A) -6x + 3y = 18 | | not same not same | | (-1)*(4x - 3y = 6) -> -4x +3y = -6 B) 3*(2x + 4y = 24) -> 6x + 12y = 72 | not same | 6x + 12y = 36 C) (-3)*(3x - y = 14) -> -9x + 3y = -42 | | | -9x + 3y = -42 D) 5x + 2y = 13 | | not same not same | | (-5)(-x + 4y = -6) -> 5x - 20y = 30